OpenStudy (anonymous):

the total area bounded by the curve y^2(1-x) = x^2(1+x) and the line x = 1 ???

OpenStudy (kamille):

Can you give us a picture?: )

OpenStudy (anonymous):

there is no picture

OpenStudy (kamille):

cant you make one by youself?

OpenStudy (anonymous):

actually i don't know how 2 solve it

OpenStudy (kamille):

Have you learnt anything about integrals?

OpenStudy (anonymous):

yes

OpenStudy (kamille):

so, firstly you need to solve eq y^2(1-x) = x^2(1+x) to see where this graphs intersect. Can you do it? Also, I add the picture how the scheme should look like.

OpenStudy (kamille):

But sadly, I dont know hot to solve it, maybe someone can help you : ) @Directrix

OpenStudy (anonymous):

k

OpenStudy (kamille):

@amistre64

OpenStudy (amistre64):

nice symmetric curves you got there; can we isolate one symmetry of it?

OpenStudy (amistre64):

and maybe define the wonky loopy part in an equivalent manner?

OpenStudy (amistre64):

ok, looks like the symetry is located at the origin

OpenStudy (amistre64):

\[ y^2(1-x) = x^2(1+x)\] \[ y^2 = x^2\frac{1+x}{1-x}\] \[ y = x\sqrt{\frac{1+x}{1-x}}\] which might be able to be integrated by parts; form x=0 to x=1

OpenStudy (anonymous):

k

OpenStudy (amistre64):

trying to solve for x is a pain :) so lets run that thru a grap and see if we get something useful

OpenStudy (anonymous):

k

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=+y%5E2%281-x%2B1%29+%3D+%28x-1%29%5E2%281%2Bx-1%29%2C+y%3Dxsqrt%28%281%2Bx%29%5C%281-x%29%29 can you see that? i shifted the original over by 1 to compare with, so I think we got a match

OpenStudy (anonymous):

k

OpenStudy (amistre64):

therefore:\[2\int_{0}^{1}x\sqrt{\frac{1+x}{1-x}}dx\] might have to do a limiting thing since at x=1 its a bad math nono \[2~\lim_{b\to 1}\int_{0}^{b}x\sqrt{\frac{1+x}{1-x}}dx\]

OpenStudy (amistre64):

if i could see a way to get an equivalent side view it might turn simpler, or even a way to go parametric

OpenStudy (anonymous):

how

OpenStudy (amistre64):

dunno yet, my brain has a mind of its own :)

OpenStudy (amistre64):

.... and of course i have to deal with bad users in the biology chat :/

OpenStudy (amistre64):

im back .... for the moment. any progress?

OpenStudy (anonymous):

i don't know how to solve further

OpenStudy (amistre64):

well, the solution is (4+pi)/2 so im sure this thing is spose to convert to polars

OpenStudy (anonymous):

k

OpenStudy (amistre64):

y^2(1-x) = x^2(1+x) r^2 = x^2 + y^2 x = rcos(t), y=rsin(t) r^2sin^2(t)(1-rcos(t)) = r^2 cos^2(t)(1+rcos(t)) sin^2(t)(1-rcos(t)) = cos^2(t)(1+rcos(t)) sin^2(t)-rcos(t)sin^2(t) = cos^2(t)+rcos^3(t) sin^2(t)-cos^2(t) = rcos^3(t) + rcos(t)sin^2(t) \[r=\frac{sin^2(t)-cos^2(t)}{cos(t)(cos^2(t)+sin^2(t))}\] \[r=\frac{sin^2(t)-cos^2(t)}{cos(t)}\]

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=polar+r%3D%28sin%5E2%28t%29-cos%5E2%28t%29%29%2F%28cos%28t%29%29 well, at least its the right shape still :)

OpenStudy (amistre64):

can we by any chance simplify that r= any more?

OpenStudy (amistre64):

sin^2 = 1 - cos^2 (1 - 2cos^2)/cos sec - 2cos

OpenStudy (amistre64):

i think thats as simple as we can get it

OpenStudy (amistre64):

be back later, have a class to get to

OpenStudy (anonymous):

k

OpenStudy (amistre64):

ok, heres what of come up with in my deranged state :)